An IQ test has a mean of 105 and a standard deviation of 20. Find the corresponding z-score for each IQ.
A.) 88
B.) 122
C.) 110
Show full solution, thank you!

Answer by math_tutor2020(3816)


I'll do part A to get you started.

The z score formula is
z = (x - mu)/sigma
where,
x = raw score
mu = mean
sigma = standard deviation

In this case,
mu = 105
sigma = 20
This applies to parts A through C

For part A, we plug in those values mentioned and x = 88 to find...
z = (x - mu)/sigma
z = (88 - 105)/20
z = -17/20
z = -0.85 is the answer to part A.

A z score of -0.85 means we're exactly 0.85 standard deviations below the mean.

I'll let you try parts B and C on your own.
Let me know if you still have questions, or if you need me to check your answers.